# Joint Learning for Scattered Point Cloud Understanding with Hierarchical Self-Distillation

### v1

@@ Abstract
Numerous point-cloud understanding techniques focus on whole entities and have succeeded in obtaining satisfactory results and limited sparsity tolerance. However, these methods are generally sensitive to incomplete point clouds that are scanned with flaws or large gaps. To address this issue, in this paper, we propose an end-to-end architecture that compensates for and identifies partial point clouds on the fly. First, we propose a cascaded solution that integrates both the upstream and downstream networks simultaneously, allowing the task-oriented downstream to identify the points generated by the completion-oriented upstream. These two streams complement each other, resulting in improved performance for both completion and downstream-dependent tasks. Second, to explicitly understand the predicted points’ pattern, we introduce hierarchical self-distillation (HSD), which can be applied to arbitrary hierarchy-based point cloud methods. HSD ensures that the deepest classifier with a larger perceptual field and longer code length provides additional regularization to intermediate ones rather than simply aggregating the multi-scale features, and therefore maximizing the mutual information between a teacher and students. We show the advantage of the self-distillation process in the hyperspaces based on the information bottleneck principle. On the classification task, our proposed method performs competitively on the synthetic dataset and achieves superior results on the challenging real-world benchmark when compared to the state-of-the-art models. Additional experiments also demonstrate the superior performance and generality of our framework on the part segmentation task.
Index Terms:
Joint Learning, hierarchical self-distillation, scattered point cloud understanding, information bottleneck.I Introduction
@@
An intuitive plan is to first compensate the incomplete parts, and then to sequentially apply a recognition algorithm on the completed point clouds. Nevertheless, models trained separately may suffer from lower robustness caused by the inconsistency in point distributions. Without explicitly learning the overall shape, as illustrated in Fig. 1(a), the network may allocate excessive attention to irrelevant points or outliers caused by irregular local shapes, leading to degraded performance. The masked autoencoder (MAE) was initially developed to pre-train encoders by learning the correspondence between masked sections and original input data such as texts, images, or point clouds. This facilitated fine-tuning processes for natural language processing (NLP) [6] and computer vision [7, 8, 9] tasks. Albeit MAE is primarily regarded as a pre-training model for improving classification performance, its reconstruction capacity is advantageous in scenarios where observations are incomplete. Additionally, hierarchy-based networks are designed to aggregate sparse features learned from multiple scales for better coverage of local regions. However, as shown in in Fig. 1(b) (left), the aggregation operation may result in the aggregated representation being unaware of the mutual information across scales. As a solution, Deeply-Supervised Nets (DSN) [10] provides separate guidance to each scale as a strong regularization technique, but is still unaware of the correlation among those scales.
@@
In order to tackle the challenges inherent in scattered point cloud classification, we propose an approach that jointly learns the completeness and shape category with hierarchical self-distillation (HSD). It consists of two cascaded streams: upstream and downstream. In the upstream, we generate a small number of portions of a point cloud based on randomly selected centroids that serve as the unmasked parts. Thereafter, a MAE is leveraged to learn the asymmetric mapping from the unmasked (complete) and masked (incomplete) point clouds. In the downstream, we can use any hierarchy-based point cloud classifier, such as PointNet++ [2], CurveNet [11], or PointMLP [12]. To further regularize the training process, our HSD framework aims to maximize the mutual information from multiple coverage scales and transfers the most distinct features in the deepest hyperspace to other scales, as shown in Fig. 1(b) (right). Specifically, the last classifier in the series from different scales is designated as the teacher to supervise the previous ones, thus transferring the important knowledge to different hyperspaces. We show that DSN can be viewed as a special case of our HSD. Theoretically, we leverage the information bottleneck principle to explain our model and demonstrate its generalization performance in experiments.
Our method places emphasis on joint learning, where both MAE and HSD are leveraged to harmonize distributions, thus enhancing the efficacy of both tasks, i.e., reconstruction and understanding. Our primary contributions are summarized as follows:
@@
A hierarchical self-distillation (HSD) framework is proposed to maximize the mutual information captured in scale-aware hyperspaces. Its plug-and-play property is highly flexible and therefore can be integrated to any hierarchy-based point cloud baselines.
Our evaluation results on scattered point clouds demonstrate that our method performs competitively on ModelNet40, while also attains state-of-the-art (SOTA) performance on the real-world dataset ScanObjectNN for classification, as well as ShpeNetPart for part segmentation.
II Related Work
Masked Autoencoder. Initially, the masked autoencoder (MAE) was introduced as a self-supervised solution to handle the lack of accessible labels in data for NLP [6]. Although the information density in computer vision is different from that of NLP, researchers have shown that MAE works effectively in computer vision by masking out certain portions of pixels in the asymmetric architecture to effectively learn largely missing patches in the pixel space [7, 8, 9, 13]. Following the image works, PointMAE [14] was the first MAE method for point clouds. Based on PointMAE, more 3D related MAEs [15, 16] have been proposed.
@@ II Related Work
Conventional Knowledge Distillation. Aiming at improving the performance of deep learning models, a common way is to average all predictions on the same data. This cumbersome scheme expedites the invention of conventional knowledge distillation that transfers the knowledge among different models [24]. Generally, teacher networks with greater model complexity learn knowledge and guide student networks with more compactness to learn from better representations for enhanced generalization. As a result, student networks are more suitable for deployment on systems with less computational resource [25, 26], e.g., mobile phone and circuit board. Other investigations into knowledge distillation, with an emphasis on early stopping of teacher’s training, have asserted that fully converged teachers do not typically promote student learning [27, 28]. In this study, we trivially bypass the need for optimal teacher selection through end-to-end self-distillation.
@@
Self-Distillation. Born-again networks [29] revisited knowledge distillation with a different objective by disentangling the training procedure from model compression. In a progressive fashion, the student models that are identical to the teacher model are trained consecutively from the supervisions of previous generations, forming an ensemble at the end that even outperforms the teacher model at the first generation.
@@
Recently, a new concept called self-distillation has emerged following the development of born-again networks. This can be divided into two categories based on whether knowledge is distilled within [30, 31, 32] or out of [33, 34] the same iteration in the training stage, for tasks such as label smoothing or weakly supervised learning perturbation. Our proposed method belongs to the former one, in which we train each classification model simultaneously regardless of the roles. Inspired by those born-again networks, we tend to transfer the knowledge from the deepest level of the hierarchy to shallower models through self-distillation.
@@
Point Cloud Understanding. Studies on point cloud have demonstrated outstanding performances on the publicly recognized benchmarks. Pioneered by PointNet [1], hierarchical scaling networks were proposed to enhance the global and local awareness, including PointNet++ [2], DGCNN [3], GBNet [35], CurveNet [11], PointMLP [12], Point-PN [36] and PointConT [37]. PointNet++ and DGCNN formulate the hierarchical structures by learning the point relation from multi-scale grouping (MSG) or k nearest neighbors (kNN) in the local region. CurveNet walks along the sequence of connected points, forming hypothetical curves that augment pointwise features. PointMLP rethinks PointNet++ with simple yet effective residual connections, holding the state of the art. In this paper, we meticulously select several well-established hierarchy-based networks (i.e., PointNet++, CurveNet, PointMLP) in our joint learning framework for the downstream task.
III Method
@@
Problem Formulation. Given the complete point set and its incomplete counterpart , we seek for a solution to reconstruct from and simultaneously classify/segment the reconstructed , where , , and denote the number of each point set in the Euclidean space, respectively. The architecture of our method is illustrated in Fig. 2. Our end-to-end architecture is composed of upstream and downstream sub-networks, denoted as and .
III-A Preliminary
@@
Strong Regularization. To solve the gradient vanishing problem, Deeply-Supervised Nets (DSN) [10] simultaneously minimizes loss functions across hidden layers [39]. DSN functions as a strong regularization and it smooths the gradient in the back propagation. From the perspective of IB principle, the optimization of DSN can be formulated as:
@@
where controls the emphasis on preserving relevant information. In a cascaded manner, each layer performs analogous to an ensemble of previous layers (companions). Obviously, the second term can be regarded as the regularization over a conventional single objective function. Intuitively, DSN treats both the output layer and companions equally.
III-B Upstream Network
@@
In real scenes, it is common that the scanned point clouds are incomplete or imperfect due to sensor resolution, obstacles, or vibration. To alleviate this phenomenon, we adopt a simplified version of SnowflakeNet to be our upstream network to learn the mapping between and , which is optimized by a reconstruction loss:
where denotes the squared Euclidean 2-norm of Chamfer distance (CD).
III-C Downstream Network
@@
Generally, the downstream network is dependent upon varying tasks. Targeting on point cloud classification problem, we leverage three elite versions of canonical hierarchy-based methods, i.e., PointNet++, CurveNet, and PointMLP, as our downstream networks, which have demonstrated satisfactory local awareness. The objective in this task obeys Eq. 2, which is cross-entropy loss, denoted as . Without losing generality, most point cloud processing methods have aggregation operations (e.g., max pooling, summation, and global set abstraction) at the end of feature extraction, because of the orderless nature of point clouds. However, such operations potentially eliminate the distinguishable features, which leads to degraded performances. To effectively learn the knowledge across different scales, we propose to use hierarchical self-distillation as a novel feature fusion framework, which will be described in the next section.
III-D Stronger Regularization
@@
Hierarchy-based networks are capable of producing multi-level features that are generated from different scales, e.g., the integrated features from MSG of PointNet++ in Fig. 1(b). However, MSG-like schemes have limitations, as they do not capture the dissociated information among features. Hence, we propose a solution that involves building multiple classification heads for hierarchical feature extraction and utilizing self-distillation to learn comprehensive knowledge among all levels. We dub this scheme as PointHSD, a general framework that can be employed by arbitrary hierarchy-based point cloud methods. A universal structure for hierarchical feature extraction (HFE) that includes the elite versions of classical point cloud classification networks (i.e., PointNet++, CurveNet, and PointMLP) is illustrated in Fig. 3. Here, set abstraction, curve convolution, and residual MLPs serve as local feature extractors that already embed local aggregation layers inside. For further details, please refer to the supplementary material.
@@
The Hierarchical Feature Extraction (HFE) module, as shown in Fig. 2, is responsible for learning the high-level local details from the low-level reconstructed point set . This is achieved by using different branches, each of which represents a distinct range of surroundings. Specifically, let denote this procedure, yielding intermediate and deepest latent representations and . We want to have a larger perceptual field, such that the difference and between and intermediate can be learned by self-distillation. The regularization term in Eq. 2 is therefore reformulated as:
By maximizing the mutual information , we can equivalently quantify the difference (denoted as ) by minimizing the Kullback-Leibler (KL) divergence. The third term in Eq. 4 further regularizes the optimization by smoothing the hard predicted labels of students (Fig. 1(b)). As such, our HSD framework treats the output layer and companions distinctively. Apparently, deep supervision [10] can be seen as a special case of Eq. 4 if .
Lemma If a teacher has a larger local perceptual field and code length than students, then .
Proof Assume for any , holds, then:
Since , we have . Along with the cross-entropy term pertaining to students, the conditional mutual information (equivalently quantified by ) imparts smoothness to the probabilistic distribution through the convolution of information derived from the teacher. This mechanism contributes to stronger regularization of the students, thereby ultimately mitigating the overconfidence of the ensemble model.
III-E Joint Learning
@@
Recall the intuitive example in Fig. 1(a), as the resolution or completeness of input point clouds declines, the accuracy of classification methods tend to decrease substantially. This is due to the fact that incomplete shapes can lead to misleading local features.
@@
In order to overcome this issue, we introduce a simple yet effective function to jointly learn the completeness and category of objects, denoted as:
@@
where means the cascade of two streams, and follows the Markov chain with the assumption that . Therefore, we designate the deepest layer as the teacher since it contains most comprehensive knowledge to smooth the hard labels in the companions. Mutual information is implicitly compressed by the network. Finally, combining Eqs. 2, 3, and 4, our joint learning objective is optimized by
where , , are the weights, which are empirically set to 1, 0.001, and 0.8, respectively.
It is important to note that our joint learning framework shares similarities with the cascaded fusion network, CasFusionNet [40], designed for scene completion and semantic segmentation. The fundamental distinction lies in that CasFusionNet focuses on dense feature fusion within the hyperspace, amalgamating both geometric and semantic features, while our method extends its influence through the integration of both low-level and high-level features to achieve joint promotion. Our work also bears similarities with GBNet [35] that leverages both low-level and high-level geometric cues. However, the substantial departure lies in that GBNet still simply aggregates local contexts, whereas our HSD perceives and distills dissociated information through features from different scales, as illustrated in Fig. 1(b). Furthermore, our joint learning design imparts a unique capability to our pipeline, making it particularly effective in handling scattered point clouds.
IV Experiment

### v3

@@ Abstract
Numerous point-cloud understanding techniques focus on whole entities and have succeeded in obtaining satisfactory results and limited sparsity tolerance. However, these methods are generally sensitive to incomplete point clouds that are scanned with flaws or large gaps. In this paper, we propose an end-to-end architecture that compensates for and identifies partial point clouds on the fly. First, we propose a cascaded solution that integrates both the upstream masked autoencoder (MAE) and downstream understanding networks simultaneously, allowing the task-oriented downstream to identify the points generated by the completion-oriented upstream. These two streams complement each other, resulting in improved performance for both completion and downstream-dependent tasks. Second, to explicitly understand the predicted points’ pattern, we introduce hierarchical self-distillation (HSD), which can be applied to any hierarchy-based point cloud methods. HSD ensures that the deepest classifier with a larger perceptual field of local kernels and longer code length provides additional regularization to intermediate ones rather than simply aggregating the multi-scale features, and therefore maximizing the mutual information (MI) between a teacher and students. The proposed HSD strategy is particularly well-suited for tasks involving scattered point clouds, wherein a singular prediction may yield imprecise outcomes due to the inherently irregular and sparse nature of the geometric shape being reconstructed. We show the advantage of the self-distillation process in the hyperspaces based on the information bottleneck principle. Our method achieves state-of-the-art on both classification and part segmentation tasks.
1 Introduction
@@
An intuitive plan is to first compensate the incomplete parts, and then to sequentially apply a recognition algorithm on the completed point clouds. Nevertheless, models trained separately may suffer from lower robustness caused by the inconsistency in point distributions. Without explicitly learning the overall shape, the network may allocate excessive attention to irrelevant points or outliers caused by irregular local shapes, leading to degraded performance. The masked autoencoder (MAE) was initially developed to pre-train encoders by learning the correspondence between masked sections and original input data such as texts, images, or point clouds. This facilitated fine-tuning processes for natural language processing (NLP) [6] and computer vision [7, 8, 9] tasks. Albeit MAE is primarily regarded as a pre-training model for improving classification performance, its reconstruction capacity is advantageous in scenarios where observations are incomplete. Additionally, hierarchy-based networks are designed to aggregate sparse features learned from multiple scales for better coverage of local regions. However, the aggregation operation may result in the aggregated representation being unaware of the mutual information across scales. As a solution, Deeply-Supervised Nets (DSN) [10] provides separate guidance to each scale as a strong regularization technique, but is still unaware of the correlation between those scales.
@@
In order to tackle the challenges inherent in scattered point cloud classification, we propose an approach based on the previous work [11] for jointly learning object completeness and shape category through hierarchical self-distillation (HSD). In addition, we extend this approach to part segmentation by developing an adapted version of HSD that enhances point-wise understanding using richer shape representations. Our network consists of two cascaded streams: upstream and downstream. In the upstream, we generate portions of a point cloud by selecting random centroids as unmasked parts, after which a MAE learns the asymmetric mapping from unmasked (complete) to masked (incomplete) point clouds. In the downstream, we can use any hierarchy-based point cloud classifier, such as PointNet++ [2], CurveNet [12], or PointMLP [13]. To further regularize training, our HSD framework maximizes mutual information across multiple coverage scales and transfers the most distinctive features from the deepest hyperspace to other scales. Specifically, the classifier at the last scale serves as the teacher to supervise preceding classifiers, thus transferring critical knowledge across different hyperspaces. We demonstrate that DSN can be viewed as a special case of HSD and leverage the information bottleneck principle to explain our model, validating its generalization performance through experiments. Our method emphasizes joint learning, where both MAE and HSD harmonize distributions to improve reconstruction and understanding. Our primary contributions are summarized as follows:
@@
We introduce a hierarchical self-distillation (HSD) framework that maximizes the mutual information captured in scale-aware hyperspaces. With its plug-and-play flexibility, HSD can be integrated into any hierarchy-based point cloud baseline.
Our evaluations on scattered point clouds show that our method performs competitively on ModelNet40, achieving state-of-the-art (SOTA) results on the real-world ScanObjectNN dataset for classification and on ShapeNetPart for part segmentation.
2 Related Work
2.1 Masked Autoencoder.
Initially, the masked autoencoder (MAE) was introduced as a self-supervised solution to handle the lack of accessible labels in data for NLP [6]. Although the information density in computer vision is different from that of NLP, researchers have shown that MAE works effectively in computer vision by masking out certain portions of pixels in the asymmetric architecture to effectively learn largely missing patches in the pixel space [7, 8, 9]. Following the image works, PointMAE [14] was the first MAE method for point clouds. Based on PointMAE, more 3D related MAEs [15, 16] have been proposed.
@@ 2 Related Work
Aiming at improving the performance of deep learning models, a common way is to average all predictions on the same data from multiple models to reduce model variance and generalization error. This cumbersome scheme expedites the invention of conventional knowledge distillation that transfers the knowledge among different models [24]. Generally, teacher networks with greater model complexity learn knowledge and guide student networks with more compactness to learn from better representations for enhanced generalization. As a result, student networks are more suitable for deployment on systems with less computational resource [25, 26], e.g., mobile phone and circuit board. Other investigations into knowledge distillation, with an emphasis on early stopping of teacher’s training, have asserted that fully converged teachers do not typically promote student learning [27, 28]. In this study, we trivially bypass the need for optimal teacher selection through end-to-end self-distillation.
2.3 Self-Distillation.
@@
Born-again networks [29] revisited knowledge distillation with a different objective by disentangling the training procedure from model compression. In a progressive fashion, the student models that are identical to the teacher model are trained consecutively from the supervisions of previous generations, forming an ensemble at the end that even outperforms the teacher model at the first generation.
@@
Recently, a new concept called self-distillation has emerged following the development of born-again networks. This can be divided into two categories based on whether knowledge is distilled within [30, 31, 32] or out of [33, 34] the same iteration in the training stage, for tasks such as label smoothing or weakly supervised learning. Our proposed method belongs to the former one, in which we train each classification model simultaneously regardless of the roles. Inspired by those born-again networks, we tend to transfer the knowledge from the deepest level of the hierarchy to shallower models through self-distillation.
2.4 Point Cloud Understanding.
@@
Studies on point cloud have demonstrated impressive performance across widely recognized benchmarks. Pioneered by PointNet [1], various hierarchical scaling networks have been proposed to enhance both global and local feature awareness, including PointNet++ [2], DGCNN [3], CurveNet [12], PRA-Net [35], MKANet [36], PointMLP [13], PointRas [37], Point-PN [38], APP-Net [39], SPoTr [40], PointConT [41], GLSNet++ [42] and PointGST [43]. Networks such as PointNet++ and DGCNN establish hierarchical structures by capturing point relationships through multi-scale grouping (MSG) or k-nearest neighbors (kNN) within local regions. CurveNet traverses sequences of connected points, creating hypothetical curves that enrich pointwise features, while PointMLP reinterprets PointNet++ with simple yet effective residual connections, achieving state-of-the-art results. In this study, we incorporate several well-established hierarchy-based networks (i.e., PointNet++, CurveNet, PointMLP) in our joint learning framework to support downstream tasks.
3 Method
3.1 Preliminary
@@
Problem Formulation. Given the complete point set and its incomplete counterpart , we seek for a solution to reconstruct from and simultaneously classify or segment the reconstructed , where and denote the number of each point cloud within Euclidean space. The architecture of our method is illustrated in Fig. 1. Our end-to-end architecture is composed of upstream and downstream sub-networks, denoted as and , respectively.
@@
Strong Regularization. To solve the gradient vanishing problem, Deeply-Supervised Nets (DSN) [10] simultaneously minimizes loss functions across recognition layers [45]. DSN acts as a strong regularization, smoothing gradients during backpropagation. From the perspective of IB principle, DSN optimization can be formulated as:
@@
where controls the emphasis on preserving relevant information. In a cascaded manner, each layer functions analogously to an ensemble of previous layers, referred to as companions. Notably, the second term serves as a regularization over a conventional single objective function. Intuitively, DSN treats both the output layer and intermediate companion layers equally.
3.2 Upstream Network
@@
In real scenes, scanned point clouds are often incomplete or imperfect due to factors such as sensor resolution limitations, obstacles, or vibrations. To alleviate this phenomenon, we adopt a simplified version of SnowflakeNet as our upstream network , which learns the mapping between and and is optimized using a reconstruction loss:
where denotes the squared Euclidean 2-norm of Chamfer distance (CD). This objective in Eq. 3 quantifies the bidirectional similarities between the two point sets, promoting the reconstructed point cloud’s proximity to the target point cloud.
3.3 Downstream Network
@@
Generally, the downstream network is dependent upon varying tasks. Targeting on point cloud classification problem, we employ three elite versions of canonical hierarchy-based methods, i.e., PointNet++, CurveNet, and PointMLP, as our downstream networks, each of which has demonstrated satisfactory local awareness. The objective for this task obeys Eq. 2, defined by the cross-entropy loss, denoted as . Without losing generality, most point cloud processing methods incorporate aggregation operations (e.g., max pooling, summation, and global set abstraction) at the end of feature extraction to accommodate the orderless nature of point clouds. However, these operations may unintentionally eliminate distinguishable features, leading to degraded performances. To address this, we propose a hierarchical self-distillation framework for feature fusion across different scales, as described in the following section.
3.4 Stronger Regularization
@@
Hierarchy-based networks generate multi-level features derived from various scales or levels, e.g., the integrated features from MSG of PointNet++ [2] and the intermediate features from multihead attentions of GSRFormer [46] or from geometry-based multiprojection fusion modules of [47]. However, MSG-like schemes have limitations, as they do not capture the dissociated information among features. Intuitively, an incomplete object might be encoded incorrectly if it lacks a critical part essential for its representation. Hence, our approach involves constructing multiple classification heads for hierarchical feature extraction, coupled with self-distillation to capture comprehensive knowledge across all levels. We term this approach PointHSD, a general framework applicable to any hierarchy-based point cloud method. A universal structure for hierarchical feature extraction (HFE) that includes the elite versions of classical point cloud classification networks (i.e., PointNet++, CurveNet, and PointMLP) is depicted in Fig. 2. Here, set abstraction, curve convolution, and residual MLPs act as local feature extractors, incorporating with local aggregation layers in each module.
@@
The Hierarchical Feature Extraction (HFE) module, as shown in Fig. 1, is responsible for learning high-level local details from the low-level reconstructed point set . This is achieved by using different branches, each representing a distinct range of surroundings, i.e., by using , , and to extract kNN features in the branches. The outputs of these branches are aligned to identical dimensions through 3 independent fully-connected layers, facilitating subsequent distillation processes. Let denote this procedure, which generates intermediate and deepest latent representations and , respectively. We expect to have a larger perceptual field, such that the difference between and the intermediate representations can be effectively captured through self-distillation. The regularization term in Eq. 2 is thus reformulated as:
where by maximizing the mutual information , we can equivalently quantify the difference (denoted as ) by minimizing the Kullback-Leibler (KL) divergence. Accordingly, the third term in Eq. 4 can be implemented as to further regularize the optimization by smoothing the hard predicted labels of students. DSN in Eq. 2 aims to enhance feature quality across all hidden layers, functioning as a type of feature regularization to alleviate overfitting [10]. During backpropagation, all companion outputs are treated equivalently, ensuring that features from different hidden layers are associated to the common proxy, rather than solely with the final layer. In contrast, the proposed HSD approach not only emphasizes the collective contributions of all companions to enhance feature quality but also assigns a supervisory role to the last layer, acting as a “teacher” to guide preceding, potentially uncertain, companions toward a more refined representation by maximizing the mutual information. This mechanism allows the last layer to provide more informative guidance to earlier layers, alleviating overfitting issues in companion layers. More intuitively, this strategic supervision aims to support the iterative refinement of the classification process, especially in scenarios where multiple attempts at classification are needed to achieve an accurate representation of a roughly reconstructed shape. Therefore, our HSD framework distinguishes between the output layer and its companions. Notably, deep supervision can be seen as a special case of Eq. 4 if .
Mathematically, as may still lack certain regions, the reconstructed shape could contain unrelated or less relevant features, such that . For any , let be the nearest neighbors for feature extraction in each layer. As increases, the mutual information between the class label and the latent feature also increases, reflecting a more accurate representation of the 3D point cloud due to the integration of broader neighborhood information. Hence, in Eq 4 facilitates the transfer of broader contextual features from layer to flow back to previous ones. Alongside the cross-entropy term associated with student layers, the term imparts smoothness to the probabilistic distribution through the convolution of information derived from the teacher. This mechanism contributes to stronger regularization of the students, thereby ultimately mitigating the overconfidence of the ensemble model.
3.5 Joint Learning
@@
Recall that as the resolution or completeness of input point clouds declines, classification accuracy tends to decline significantly, primarily because incomplete shapes can yield misleading local features.
@@
In order to overcome this issue, we introduce a straightforward yet effective function to jointly learn both object completeness and category, defined as:
@@
where denotes the cascade of two streams, and follows the Markov chain with the assumption that . Therefore, we designate the deepest layer as the teacher, as it encapsulates most comprehensive knowledge to soften hard labels in the other layers. The mutual information is implicitly compressed by the network. Finally, combining Eqs. 2, 3, and 4, we optimize our joint learning objective as follows:
where , , are the weights, which are empirically set to 1, 0.001, and 0.8, respectively. and control the balance between upstream and downstream processes, which are guided by the sparse data-focused paradigm design that requires reconstruction as a prerequisite. Additionally, distorts the amount of information used for direct optimization in classification or for flowing back to regularize the network. In this way, our network learns to sequentially and simultaneously reconstruct and recognize scattered point clouds, merging both geometric and semantic priors into the model.
It is important to note that our joint learning framework shares similarities with the cascaded fusion network, CasFusionNet [48], which is designed for scene completion and semantic segmentation. However, a fundamental distinction lies in that CasFusionNet focuses on dense feature fusion within the hyperspace, amalgamating both geometric and semantic features, while our method extends its influence by integrating both low-level and high-level features to achieve a synergistic enhancement across tasks. Moreover, unlike CasFusionNet, the plug-and-play property of our framework allows either the upstream or downstream network to be replaced to better suit specific application needs.
4 Experiment